OpenStudy (puertoricangirl):

The table below shows the number of marbles of different colors in a bag: .Color of Marbles .Number of Marbles Pink 5 Blue 2 Green 3 Ursula draws a marble from the bag randomly without looking. She then draws another marble from the bag without replacing the first one. Which expression shows the probability of drawing green marbles in both the trials? A. 3 over 10 multiplied by 2 over 9 B.3 over 10 multiplied by 2 over 10 C.3 over 10 added to 2 over 9 D. 3 over 10 added to 2 over 10

jimthompson5910 (jim_thompson5910):

There are 5+2+3 = 10 marbles total The probability of picking green is 3/10 because there are 3 green out of 10 total We do NOT replace the marble we picked, so we have 10-1 = 9 marbles left. There would be 3-1 = 2 green marbles left On the second pick, the chances of getting green is 2/9 because there are 2 green out of 9 At this point, we multiply out the probabilities (3/10) * (2/9)

jimthompson5910 (jim_thompson5910):

most books would want you to multiply out (3/10) * (2/9) but it appears your book wants you to stop here

OpenStudy (puertoricangirl):

any one else wants to say what they think? the answer

OpenStudy (puertoricangirl):

@niuuu since you r there why dont you tell me opinion ?

OpenStudy (puertoricangirl):

@Luigi0210

OpenStudy (robtobey2):

A.

OpenStudy (puertoricangirl):

ok thx !